A particle is moving with the given data. Find the position of the particle.
a(t) = t - 2
s(0) = 1
v(0) = 1

s(t) =_________
1
Expert's answer
2011-03-17T05:58:38-0400
v = int(a(t) dt) = t2/2 - 2t +v(0) = t2/2 - 2t + 1
s = int(v(t) dt) = t3/6 - t2 + t +s(0) = t3/6 - t2 + t + 1